# Revealing the Inherent Instructability of Pre-Trained Language Models

### v2

@@ 2. Refinement: Edit the extracted text to enhance readability and presentation:
Table 24: Example of a refined LIMA response. The response is refined using Llama-3.1-70B-Instruct with the refinement prompt in Table 23. We observe the refined response exhibits improved coherence.
@@
Llama-3.1-8B (Dubey et al., 2024) IT Alpaca (response refined) 53.37
@@
Gemma-2-9B (Riviere et al., 2024) IT Alpaca (response refined) 56.30
@@
Mistral-7B-v0.3 (Jiang et al., 2023) IT Alpaca (response refined) 56.02
@@
Gemma-2-2B (Riviere et al., 2024) IT Alpaca (response refined) 52.16
@@
Table 25: GPT-4 preference evaluation results for IT and RT models trained with refined responses. These models largely outperform their counterparts trained on original responses.
@@
Table 27: Evaluation results for open-data OLMo IT/RT models. We find results similar to those in Sections 4 and 6.

### v3

@@ 2. Refinement: Edit the extracted text to enhance readability and presentation:
Table 28: Example of a refined LIMA response. The response is refined using Llama-3.1-70B-Instruct with the refinement prompt in Table 27. We observe that the refined response exhibits improved coherence.
@@
Llama-3.1-8B
(Dubey et al., 2024)
IT Alpaca (response refined) 53.37
@@
Gemma-2-9B
(Riviere et al., 2024)
IT Alpaca (response refined) 56.30
@@
Mistral-7B-v0.3
(Jiang et al., 2023)
IT Alpaca (response refined) 56.02
@@
Gemma-2-2B
(Riviere et al., 2024)
IT Alpaca (response refined) 52.16
@@
Table 29: GPT-4 preference evaluation results for RT and IT models trained with refined responses. These models largely outperform their counterparts trained on original responses.
@@
Table 31: Evaluation results for open-data OLMo RT/IT models. We find results similar to those in Section 4 and 6.